# Well-balanced high order finite difference WENO schemes for a first-order Z4 formulation of the Einstein field equations

### v1

@@ Abstract
In this work we aim at developing a new class of high order accurate well-balanced finite difference (FD) Weighted Essentially Non-Oscillatory (WENO) methods for numerical general relativity, which can be applied to any first–order reduction of the Einstein field equations, even if non–conservative terms are present. We choose the first–order non–conservative Z4 formulation of the Einstein equations, which has a built–in cleaning procedure that accounts for the Einstein constraints and that has already shown its ability in keeping stationary solutions stable over long timescales. Upon the introduction of auxiliary variables, the vacuum Einstein equations in first order form constitute a PDE system of 54 equations that is naturally non-conservative. While prior versions of FD WENO schemes were developed exclusively for hyperbolic systems of conservation laws, we show how to design FD WENO schemes that can handle non-conservative products. Different variants of FD–WENO are discussed, with an eye to their suitability for higher order accurate formulations for numerical general relativity. We successfully solve a set of fundamental tests of numerical general relativity with an order of accuracy that has been implemented up to ninth order in space. Due to their intrinsic robustness, flexibility and ease of implementation, finite difference WENO schemes can effectively replace traditional central finite differencing in any first–order formulation of the Einstein field equations, without any artificial viscosity. When used in combination with well balancing, the new numerical schemes allow to preserve stationary equilibrium solutions of the Einstein equations exactly, such as isolated rotating and non-rotating black holes. This is particularly relevant in view of the numerical study of the quasi-normal modes of oscillations of relevant astrophysical sources. In conclusion, general relativistic high energy astrophysics could benefit from this new class of numerical schemes and the ecosystem of other desirable capabilities that has been built around them.
@@ 1 Introduction
As is well known, the Einstein equations present inbuilt challenges that do not show up in many of the other governing equations commonly considered in computational astrophysics. For one thing, the system naturally arises as a second order PDE system, whereas our best numerical tools have been developed for first order PDE systems. Possibly due to the great results made possible by the second–order implementation of the BSSNOK formulation of the Einstein equations (Shibata & Nakamura, 1995; Baumgarte & Shapiro, 2010), most of the first–order analysis have instead remained at an exploratory or theoretical level (Fischer & Marsden, 1972; Anderson & York, 1999; Kidder et al., 2001; Bona et al., 2003; Alvi, 2002; Alekseenko & Arnold, 2003; Buchman & Bardeen, 2003; Bona & Palenzuela-Luque, 2005; Olivares et al., 2022). Notable but rare examples of practical implementations of first order formulations into numerical codes were provided by Bona et al. (2004a), Brown et al. (2012), Buchman & Bardeen (2005), Dumbser et al. (2018). The latter, in particular, applied Discontinuous Galerkin (DG) schemes to the conformal and covariant Z4 formulation of the Einstein equations (FO-CCZ4) of Alic et al. (2012). More recently, Dumbser et al. (2024) have obtained very promising results after using DG methods in a well–balanced first–order implementation of the damped version of the Z4 formulation proposed by Gundlach et al. (2005), henceforth referred to as the FO-Z4 formulation. We recall that the Z4 formulation of the Einstein equations was originally proposed by Bona et al. (2003, 2004a) to automatically account for a proper treatment of the Einstein constraints through the addition of a four vector $z^{\mu}$, in a rather similar way to what is done in the divergence cleaning approach by Munz et al. (2000) and Dedner et al. (2002) for the Maxwell and magnetohydrodynamics (MHD) equations. Later on, this approach has been combined with the conformal decomposition of the metric, which is missing in the original Z4 formulation, to obtain the so–called CCZ4 and Z4c formulations (Alic et al., 2012; Hilditch & Richter, 2012; Hilditch et al., 2012; Alic et al., 2013; Dumbser et al., 2018; Peterson et al., 2023).
@@ 2 The damped first–order Z4 system of the Einstein equations
where $G_{\mu\nu}$ is the Einstein tensor and $n^{\mu}$ is the four–velocity of the Eulerian observer. The four vector $z^{\mu}=\Theta n^{\mu}+Z^{\mu}$ is artificially introduced to clean the violations of the Einstein constraints, while the two constant coefficients $\kappa_{1}$ and $\kappa_{2}$ can act as damping mechanisms over the four vector $z^{\mu}$. Not present in the original formulation by Bona et al. (2003, 2004b), such coefficients were introduced by Gundlach et al. (2005) and slightly modified by Dumbser et al. (2024) in such a way to avoid their mutual multiplication in the resulting PDE system. The extrinsic curvature of the hypersurface $\Sigma_{t}$, a crucial quantity of differential geometry, is one of the primary variables of the Z4 formulation and it is defined as
@@
where $\mathbf{U}$ is the state vector, composed of 54 dynamical variables, i.e. 10 for the lapse, the shift vector and the metric components, 6 for $K_{ij}$, 4 for the $z^{\mu}$ four vector, 3 for $A_{i}$, 9 for $B_{i}^{\,\,j}$, 18 for $D_{ijk}$, 1 for $K_{0}$ and 3 for $b^{i}$. The source term $\bf S(\mathbf{U})$ contains algebraic terms only. The hyperbolic nature of (16) has been proved by Dumbser et al. (2024) after a careful analysis of the matrix $\mathbf{A}(\mathbf{U})$ via mathematical software packages, and the interested reader is pointed to that work, and especially the Appendix in that work, for further details. We just recall here that, as already noticed by Dumbser et al. (2018), hyperbolicity is helped by moving any spatial derivatives of the metric terms $\alpha$, $\beta^{i}$ and ${\gamma}_{ij}$ to the right hand side via the auxiliary variables (4).
@@ 3.2 FD-WENO Schemes for Hyperbolic Systems with Non-Conservative Products
In the previous sub-section we have mentioned that it is advantageous to carry out the reconstruction in the characteristic variables. In a tour de force, the 54 eigenvectors of FO-Z4 system for the full Einstein–Euler system have been derived in Dumbser et al. (2024). However, it would be prohibitive to project the entire system into that eigenspace for carrying out a reconstruction of the solution vector. Fortunately, the eigenspace splits into a 5 component subspace for the hydrodynamic system and an eigenspace for the rest of the system. Since shocks form only in the hydrodynamic sector, while in the spacetime part all the corresponding fields are linearly degenerate and no shocks can form, we treat the reconstruction of the variables in the GR sector without any characteristic projection. This is not a fundamental limitation, it is only a time-saving expedient in light of the fact that the first order hyperbolic formulation of GR results in such a large system.
It is also useful to observe that eqn. (31) is not in exact flux conservative form. An exact flux conservative form is essential (because of the Lax-Wendroff theorem) for accurately predicting shock locations when the system is conservative, as is the case for hydrodynamics. The modern trend is to seek out schemes that revert to a conservation form when such is present. However, eqn. (31) does derive from eqn. (25) which is in conservation form, so for most practical cases it does a rather good job in predicting shock locations. Besides, since the solutions of the Einstein equations are smooth, the lack of a conservation form is not a major impediment. As a result, eqn. (31) has an important place in the solution of the Einstein equations. This completes our description of FD-WENO schemes as they are extended to hyperbolic PDEs, like the FO-Z4 system, that are not in conservation form.
@@ 4 Numerical tests
In the following, we show the performances of the new schemes over a representative sample of standard tests in NR. While doing so, we are also going to monitor the energy constraint $H$ and the momentum constraints $M_{i}$, also known as ADM constraints, which are defined as
@@ 4.1 Gauge wave
Figure 4, on the contrary, reports the results obtained with FD–WENO schemes, by showing, in the first column, the profiles of the lapse $\alpha$ in a long run with a final time $t=1000$, compared to the exact solution, for various orders, and using a wave amplitude $A=0.1$. In the right column we display the evolution of the ADM constraints, which remain perfectly under control all along the simulation. Figure 5 serves the same purpose, but this time for the results obtained with AFD–WENO schemes. It is interesting to note that on such a long simulation both FD–WENO-3 and AFD–WENO-2 schemes fail to reproduce the correct solution (top–left panels of Fig. 4 and Fig. 5), with a slightly better performance of FD–WENO-3 with respect to AFD–WENO-2. For the higher-order FD–WENO and AFD–WENO schemes ($k=5,7,9$), the numerical solutions match perfectly well the analytic ones.
Figure 4: Gauge wave: The left panel (Figs. 4a, 4c, 4e, 4g) shows profiles of the lapse $\alpha$ at the final time $t=1000$ for various order accurate FD–WENO schemes. The right panel (Figs. 4b, 4d, 4f, 4h) shows the $L_{2}$-error evolution of the ADM constraints for the corresponding order scheme.
Figure 5: Gauge wave: The left panel (Figs. 5a, 5c, 5e, 5g) shows profiles of the lapse $\alpha$ at the final time $t=1000$ for various order accurate AFD–WENO schemes. The right panel (Figs. 5b, 5d, 5f, 5h) shows the $L_{2}$-error evolution of the ADM constraints for the corresponding order scheme.
@@ 4.2 Robust stability test
The so–called robust stability test is used to provide an empiric indication of the hyperbolicity of the PDE system, and it was introduced when there were very few theoretical analysis of the second order BSSNOK formulation (Beyer & Sarbach, 2004). In view of the huge progress witnessed by NR in the last twenty years, and especially because strong hyperbolicity of the present first–order Z4 formulation has been verified by Dumbser et al. (2024), this test has lost part of its relevance. Nevertheless, it is still quite useful to exclude possible exponential unstable modes which could be in principle triggered by a given numerical algorithm (Calabrese et al., 2006). We should also comment on the fact that, to the best of our knowledge, none of the first–order formulations of the Einstein equations have shown any problem with this test (Bona et al., 2004a; Brown et al., 2012; Dumbser et al., 2018, 2024). We have therefore prepared the usual setup, given by a flat Minkowski spacetime in two space dimensions, in which all the metric terms are perturbed with a random perturbation of amplitude ±$\pm 10^{-7}/\varrho^{2}$. The integer $\rho$ is also used to control the resolution in a set of increasingly refined meshes. The computational domain is given by the square $[-0.5;0.5]\times[-0.5;0.5]$. As customary for this test, the gamma-driver is activated, with $\eta=0.2$ in Eq. (7). The time evolution of the Einstein constraints for the FD-WENO and for the AFD-WENO schemes are shown in Fig. 6 and Fig. 7, respectively, establishing quite a successful outcome for this test.
@@ 4.3 Gowdy wave
We have solved this test as a one-dimensional problem using a mesh composed of 120 zones, up to the final time $\tau=1000$. We impose periodic boundary conditions. In Fig. 8 and Fig. 9 we show the results obtained with our FD-WENO and AFD-WENO schemes, respectively. The left panels report the variable $K_{xx}$, and they show a perfect matching to the exact solution at the final time. The right panels, on the other hand, show the evolution of the Einstein constraints, which, in spite of a linear increase already documented in the literature (Bona et al., 2004a), remain under control.
Figure 8: Gowdy wave: The left panel (Figs. 8a,8c,8e,8g) shows profiles of the $K_{xx}$ variable at the final time $\tau=1000$ for various order accurate FD–WENO schemes. The right panel (Figs. 8b,8d,8f,8h) shows the $L_{2}$-error evolution of the ADM constraints at the corresponding order.
Figure 9: Gowdy wave: The left panel (Figs. 9a,9c,9e,9g) shows profiles of the $K_{xx}$ variable at the final time $\tau=1000$ for various order accurate AFD–WENO schemes. The right panel (Figs. 9b,9d,9f,9h) shows the $L_{2}$-error evolution of the ADM constraints at the corresponding order.
@@ 4.4 Stationary black holes
Contrary to naive expectations, keeping a simple Schwarzschild or Kerr black hole stationary in three dimensions via a NR code is far from trivial. To the best of our knowledge, this has never been obtained through the second–order BSSNOK formulation, apart from a promising attempt performed by Alcubierre & Brügmann (2001), who, however, performed the calculations on one octant only (because of spherical symmetry) while they reported the appearance of an unstable mode when they extended the simulation in the full grid space. In 2001, after using an alternative first–order formulation of the Einstein equations, henceforth the KST formulation, Kidder et al. (2001) were able to evolve a Schwarzschild black hole in three space dimensions, though the momentum constraints did not seem to be fully under control. Better results were obtained by Brügmann (2013), who kept a Schwarzschild black hole in spherical Kerr–Schild coordinates stationary for very long timescales, using spectral methods in combination with a first–order formulation of the Einstein equations in the generalized harmonic gauge. Again in this formulation, Tichy et al. (2023) obtained rather good results as well for a non-rotating black hole, but using DG methods. Finally, after using a well–balanced version of the first–order Z4 formulation of the Einstein equations, in combination with an ADER–DG scheme, Dumbser et al. (2024) were able to preserve perfect stationary equilibria of even extreme Kerr black holes over arbitrarily long timescales.
@@
We have therefore considered the stationary solution represented by the Kerr spacetime in pseudo-Cartesian Kerr–Schild coordinates, as they can be found, for instance, in de Felice & Clarke (1990) (Sect. 11.4) or in Wiltshire et al. (2009). Here, we reconsider that problem by adopting a seventh-order FD-WENO scheme and a sixth-order AFD-WENO scheme, applied in three-dimensional (spatial) simulations. The three-dimensional computational domain is given by $\Omega=[-5;5]^{3}$ covered by a $80\times 80\times 80$ uniform grid, while, as boundary conditions, we impose the stationary equilibrium solution at the outer border. We also recall that, in the coordinates adopted, the physical singularity is mapped into the the so–called ring singularity in z = 0 plane, with a radius given by the spin of the black hole. For this reason we have excised the ring singularity from the numerical domain, which remains inside an appropriate excision box.
@@
Moreover, we have found crucial for a successful evolution of these equilibrium tests over arbitrarily long times to activate the well balancing property described in Sect. 3.5. Without well balancing, small but ever growing deviations from equilibrium where detected on times $t\geq 100$, even at the sixth order of convergence of our numerical schemes. Under these settings, we have considered the following two representative cases:
@@
  • A Schwarzschild black hole ($a=0$). Fig. 10 and Fig. 11 show the results of a long run with $t_{f}=1000\,M$, obtained with the FD-WENO scheme and with the AFD-WENO scheme, respectively. The central top panel reports a contour-color representation of the lapse $\alpha$ through the $z=0$ plane, while the arrows visualize the vector field of the shift $\beta^{i}$, which, in spite of the black hole being non/rotating, is non zero due to the choice of the coordinates. The remaining panels, with 1D profiles along the $z$ axis, show the comparison with the equilibrium solution and prove the ability of the code in preserving the stationarity of the solution.

### v2

@@ Abstract
We develop a new class of high–order accurate well-balanced finite difference (FD) Weighted Essentially Non-Oscillatory (WENO) methods for numerical general relativity, which can be applied to any first–order reduction of the Einstein field equations, even if non–conservative terms are present. We choose the first–order non–conservative Z4 formulation of the Einstein equations, which has a built–in cleaning procedure that accounts for the Einstein constraints and that has already shown its ability in keeping stationary solutions stable over long timescales. By introducing auxiliary variables, the vacuum Einstein equations in first–order form constitute a PDE system of 54 equations that is naturally non-conservative. We show how to design FD–WENO schemes that can handle non-conservative products. Different variants of FD–WENO are discussed, with an eye to their suitability for higher order accurate formulations for numerical general relativity. We successfully solve a set of fundamental tests of numerical general relativity with up to ninth order spatial accuracy. Due to their intrinsic robustness, flexibility and ease of implementation, finite difference WENO schemes can effectively replace traditional central finite differencing in any first–order formulation of the Einstein field equations, without any artificial viscosity. When used in combination with well–balancing, the new numerical schemes preserve stationary equilibrium solutions of the Einstein equations exactly. This is particularly relevant in view of the numerical study of the quasi-normal modes of oscillations of relevant astrophysical sources. In conclusion, general relativistic high–energy astrophysics could benefit from this new class of numerical schemes and the ecosystem of desirable capabilities built around them.
@@ 1 Introduction
As is well known, the Einstein equations present inbuilt challenges that do not show up in many of the other governing equations commonly considered in computational astrophysics. For one thing, the system naturally arises as a second order PDE system, whereas our best numerical tools have been developed for first order PDE systems. In this respect, it is interesting to note that, while the generalized harmonic formulation exists as a first order system since quite a while (Lindblom et al., 2006), having already allowed for highly relevant scientific results (see, among others, Boyle et al. (2007); Duez et al. (2008); Scheel et al. (2009); Szilágyi (2014); Tichy et al. (2023); Deppe et al. (2024) and references therein), the whole family represented by the BSSNOK/Z4/CCZ4 formulations (Nakamura et al., 1987; Shibata & Nakamura, 1995; Baumgarte & Shapiro, 1998; Alcubierre, 2008; Baumgarte & Shapiro, 2010; Bona & Palenzuela-Luque, 2005; Alic et al., 2009) seems to have shown a stronger inertia in performing the transition from second–order to first–order implementations, with notable but rare progresses reported by Bona et al. (2004a), Brown et al. (2012), Dumbser et al. (2018). The latter, in particular, applied Discontinuous Galerkin (DG) schemes to the conformal and covariant Z4 formulation of the Einstein equations (FO-CCZ4) of Alic et al. (2012). More recently, Dumbser et al. (2024) have obtained promising results after using DG methods in a well–balanced first–order implementation of the damped version of the Z4 formulation proposed by Gundlach et al. (2005), henceforth referred to as the FO-Z4 formulation. We recall that the Z4 formulation of the Einstein equations was originally proposed by Bona et al. (2003, 2004a) to automatically account for a proper treatment of the Einstein constraints through the addition of a four vector $z^{\mu}$, in a rather similar way to what is done in the divergence cleaning approach by Munz et al. (2000) and Dedner et al. (2002) for the Maxwell and magnetohydrodynamics (MHD) equations. Later on, this approach has been combined with the conformal decomposition of the metric, which is missing in the original Z4 formulation, to obtain the so–called CCZ4 and Z4c formulations (Alic et al., 2012; Hilditch & Richter, 2012; Hilditch et al., 2012; Alic et al., 2013; Dumbser et al., 2018; Peterson et al., 2023).
@@ 2 The damped first–order Z4 system of the Einstein equations
where $G_{\mu\nu}$ is the Einstein tensor and $n^{\mu}$ is the four–velocity of the Eulerian observer. The four vector $z^{\mu}=\Theta n^{\mu}+Z^{\mu}$ is artificially introduced to clean the violations of the Einstein constraints. We therefore notice that $\Theta$ and $Z^{\mu}$ are the 3+1 representations of $z^{\mu}$. The two constant coefficients $\kappa_{1}$ and $\kappa_{2}$ can act as damping mechanisms over the four vector $z^{\mu}$. Not present in the original formulation by Bona et al. (2003, 2004b), such coefficients were introduced by Gundlach et al. (2005) and slightly modified by Dumbser et al. (2024) in such a way to avoid their mutual multiplication in the resulting PDE system. The extrinsic curvature of the hypersurface $\Sigma_{t}$, a crucial quantity of differential geometry, is one of the primary variables of the Z4 formulation and it is defined as
@@
where $\mathbf{U}$ is the state vector, composed of 54 dynamical variables, i.e. 10 for the lapse, the shift vector and the metric components, 6 for $K_{ij}$, 3 for the three vector $Z_{i}$, 1 for the scalar $\Theta$, 3 for $A_{i}$, 9 for $B_{i}^{\,\,j}$, 18 for $D_{ijk}$, 1 for $K_{0}$ and 3 for $b^{i}$. The source term $\bf S(\mathbf{U})$ contains algebraic terms only. The hyperbolic nature of (16) has been proved by Dumbser et al. (2024) after a careful analysis of the matrix $\mathbf{A}(\mathbf{U})$ via mathematical software packages, and the interested reader is pointed to that work, and especially the Appendix in that work, for further details. We just recall here that, as already noticed by Dumbser et al. (2018), hyperbolicity is helped by moving any spatial derivatives of the metric terms $\alpha$, $\beta^{i}$ and ${\gamma}_{ij}$ to the right hand side via the auxiliary variables (4).
@@
  • The factor $\mu$ in Eq. (14) is used to insert a curl-free term that is zero on the continuum level, but that favours hyperbolicity of the whole PDE system.
@@ 3.2 FD-WENO Schemes for Hyperbolic Systems with Non-Conservative Products
In the previous sub-section we have mentioned that it is advantageous to carry out the reconstruction in the characteristic variables. In a tour de force, the 54 eigenvectors of FO-Z4 system for the full Einstein–Euler system have been derived in Dumbser et al. (2024). However, it would be prohibitive to project the entire system into that eigenspace for carrying out a reconstruction of the solution vector. Fortunately, the eigenspace splits into a 5 component subspace for the hydrodynamic system and an eigenspace for the rest of the system. We recall that, although no physical shocks can form in the metric, since all the spacetime fields are linearly degenerate, gauge shocks can still form, as a result of the gauge conditions chosen (Alcubierre, 2003; Jiménez-Vázquez & Alcubierre, 2022; Baumgarte et al., 2023b, a). In this respect, it could be quite beneficial to treat the reconstruction of the variables in the GR sector using characteristic projection. However, due to the complexity of the equations, we have not followed this approach here and we postpone the analysis of such a specialized aspect to a future investigation.
It is also useful to observe that eqn. (31) is not in exact flux conservative form. An exact flux conservative form is essential (because of the Lax-Wendroff theorem) for accurately predicting shock locations when the system is conservative, as is the case for hydrodynamics. The modern trend is to seek out schemes that revert to a conservation form when such is present. However, eqn. (31) does derive from eqn. (25) which is in conservation form, so for most practical cases it does a rather good job in predicting shock locations. Besides, since the solutions of the Einstein equations are continuous in the metric terms, except of course when a singularity forms or when gauge shocks are present, the lack of a conservation form is not a major impediment. As a result, eqn. (31) has an important place in the solution of the Einstein equations. This completes our description of FD-WENO schemes as they are extended to hyperbolic PDEs, like the FO-Z4 system, that are not in conservation form.
@@ 4 Numerical tests
In the following, we show the performances of the new schemes over a representative sample of standard tests in NR. While doing so, we are also going to monitor the energy constraint $H$ and the momentum constraints $M_{i}$, also known as ADM constraints, which, in a vacuum spacetime, are defined as
@@ 4.1 Gauge wave
Figure 4, on the contrary, reports the results obtained with FD–WENO schemes, by showing, in the first column, the numerical errors of the lapse $\alpha$ with respect to the exact solution at time $t=1000$, for various orders, using a wave amplitude $A=0.1$ and $128$ zones. In the right column we display the evolution of the ADM constraints, which remain perfectly under control all along the simulation. Figure 5 serves the same purpose, but this time for the results obtained with AFD–WENO schemes. We note that in such a long simulation both FD–WENO-3 and AFD–WENO-2 schemes lead to relatively large errors at the given resolution, as it is visible from the numbers in the top–left panels of Fig. 4 and Fig. 5, with a slightly better performance of FD–WENO-3 with respect to AFD–WENO-2. For the higher-order FD–WENO and AFD–WENO schemes ($k=5,7,9$), the numerical solutions reproduce the analytic ones with very good accuracy.
In view of possible studies of critical phenomena in gravitational collapse, we have also evolved the same gauge wave but with the 1+log gauge condition. This gauge condition is known to produce gauge shocks (Alcubierre, 1997; Baumgarte & Hilditch, 2022; Baumgarte et al., 2023a) which can easily make the code crash. The results of our computations are shown in Fig. 6, where strong gradients are formed in the lapse, but the numerical scheme can still cope with them and the code does not crash, not even after very long times. Even if at the level of an academic case, this test shows the ability of the Z4 formulation, combined with our numerical scheme, to resolve the propagation of gauge shocks. Because the shocks are only of modest strength, interpolation on the primal variables was sufficient; and interpolation in characteristic space was not needed.
Figure 4: Gauge wave: The left panel (Figs. 4a, 4c, 4e, 4g) shows the difference among the numerical solution and the exact one for the variable $\alpha$ (lapse) at time $t=1000$ for various order accurate FD–WENO schemes. The right panel (Figs. 4b, 4d, 4f, 4h) shows the $L_{2}$-error evolution of the ADM constraints for the corresponding order scheme.
Figure 5: Gauge wave: The left panel (Figs. 5a, 5c, 5e, 5g) shows the difference among the numerical solution and the exact one for the variable $\alpha$ (lapse) at time $t=1000$ for various order accurate FD–WENO schemes. The right panel (Figs. 5b, 5d, 5f, 5h) shows the $L_{2}$-error evolution of the ADM constraints for the corresponding order scheme.
Figure 6: Gauge wave: Time evolution of the lapse profile $\alpha$ at different time levels when 1+log gauge condition is being used.
@@ 4.2 Robust stability test
The so–called robust stability test is used to provide an empiric indication of the hyperbolicity of the PDE system, and it was introduced when there were very few theoretical analysis of the second order BSSNOK formulation (Beyer & Sarbach, 2004). In view of the huge progress witnessed by NR in the last twenty years, and especially because strong hyperbolicity of the present first–order Z4 formulation has been verified by Dumbser et al. (2024), this test has lost part of its relevance. Nevertheless, it is still quite useful to exclude possible exponential unstable modes which could be in principle triggered by a given numerical algorithm (Calabrese et al., 2006). We should also comment on the fact that, to the best of our knowledge, none of the first–order formulations of the Einstein equations have shown any problem with this test (Bona et al., 2004a; Brown et al., 2012; Dumbser et al., 2018, 2024). We have therefore prepared the usual setup, given by a flat Minkowski spacetime in two space dimensions, in which all the metric terms are perturbed with a random perturbation of amplitude ±$\pm 10^{-7}/\varrho^{2}$. The integer $\rho$ is also used to control the resolution in a set of increasingly refined meshes. The computational domain is given by the square $[-0.5;0.5]\times[-0.5;0.5]$, with periodic boundary conditions along each direction. As customary for this test, the gamma-driver is activated, with $\eta=0.2$ in Eq. (7). The time evolution of the Einstein constraints for the FD-WENO and for the AFD-WENO schemes are shown in Fig. 7 and Fig. 8, respectively, establishing quite a successful outcome for this test.
@@ 4.3 Gowdy wave
We have solved this test as a one-dimensional problem using a mesh composed of 120 zones, up to the final time $\tau=1000$. We impose periodic boundary conditions. In Fig. 9 and Fig. 10 we show the results obtained with our FD-WENO and AFD-WENO schemes, respectively. The left panels report the numerical error of the variable $K_{xx}$ at the final time. The right panels, on the other hand, show the evolution of the Einstein constraints, which, in spite of a linear increase already documented in the literature (Bona et al., 2004a), remain under control.
Figure 9: Gowdy wave: The left panel (Figs. 9a,9c,9e,9g) shows the difference among the numerical solution and the exact one for the variable $K_{xx}$ at time $\tau=1000$ for various order accurate FD–WENO schemes. The right panel (Figs. 9b,9d,9f,9h) shows the $L_{2}$-error evolution of the ADM constraints at the corresponding order.
Figure 10: Gowdy wave: The left panel (Figs. 10a,10c,10e,10g) shows the difference among the numerical solution and the exact one for the variable $K_{xx}$ at time $\tau=1000$ for various order accurate AFD–WENO schemes. The right panel (Figs. 10b,10d,10f,10h) shows the $L_{2}$-error evolution of the ADM constraints at the corresponding order.
@@ 4.4 Stationary black holes
Contrary to naive expectations, keeping a simple Schwarzschild or Kerr black hole stationary in three dimensions via a NR code is far from trivial. In this subsection we refer to the stationary BH solutions of Schwarzschild (1916) and Kerr (1963), both of them in horizon penetrating coordinates, and not to puncture black holes.. Within the second order BSSNOK formulation, promising attempts were performed by Alcubierre & Brügmann (2001), who, however, performed the calculations on one octant only (because of spherical symmetry) while they reported the appearance of an unstable mode when they extended the simulation in the full grid space. A substantial progress is reported in Tichy (2009), with the Hamiltonian constraint that, however, manifests an exponential growth on long timescales. In 2001, after using an alternative first–order formulation of the Einstein equations, henceforth the KST formulation, Kidder et al. (2001) were able to evolve a Schwarzschild black hole in three space dimensions, though the momentum constraints did not seem to be fully under control. Better results were obtained by Brügmann (2013), who kept a Schwarzschild black hole in spherical Kerr–Schild coordinates stationary for very long timescales, using spectral methods in combination with a first–order formulation of the Einstein equations in the generalized harmonic gauge. Again in this formulation, Tichy et al. (2023) obtained rather good results as well for a non-rotating black hole, but using DG methods. Finally, after using a well–balanced version of the first–order Z4 formulation of the Einstein equations, in combination with an ADER–DG scheme, Dumbser et al. (2024) were able to preserve perfect stationary equilibria of even extreme Kerr black holes over arbitrarily long timescales.
@@
We have therefore considered the stationary solution represented by the Kerr spacetime in pseudo-Cartesian Kerr–Schild coordinates, as they can be found, for instance, in de Felice & Clarke (1990) (Sect. 11.4) or in Wiltshire et al. (2009). Here, we reconsider that problem by adopting a seventh-order FD-WENO scheme and a sixth-order AFD-WENO scheme, applied in three-dimensional (spatial) simulations. The three-dimensional computational domain is given by $\Omega=[-5;5]^{3}$ covered by a $80\times 80\times 80$ uniform grid, while, as boundary conditions, we impose the stationary equilibrium solution at the outer border. The gauge condition adopted is the 1+log gauge condition. We also recall that, in the coordinates adopted, the physical singularity is mapped into the the so–called ring singularity in z = 0 plane, with a radius given by the spin of the black hole. For this reason we have excised the ring singularity from the numerical domain, which remains inside an appropriate excision box. Moreover, we insert an initial perturbation in the cleaning variable $\Theta$, which is chosen as
$\Theta(t=0)=A_{0}\exp\left(-\frac{1}{2}\frac{(x-2.5)^{2}+(y-0)^{2}+(z-0)^{2}}{\sigma^{2}}\right)\,,$ (58)
@@
with $A_{0}=10^{-3}$, $\sigma=0.2$. We have found crucial for a successful evolution of these equilibrium tests over arbitrarily long times to activate the well balancing property described in Sect. 3.5. Without well balancing, small but ever growing deviations from equilibrium where detected on times $t\geq 100$, even at the sixth order of convergence of our numerical schemes. Under these settings, we have considered the following two representative cases:
@@
  • A Schwarzschild black hole ($a=0$). Fig. 11 and Fig. 12 show the results of a long run with $t_{f}=1000\,M$, obtained with the FD-WENO scheme and with the AFD-WENO scheme, respectively. The central top panel reports a contour-color representation of the lapse $\alpha$ through the $z=0$ plane, while the arrows visualize the vector field of the shift $\beta^{i}$, which, in spite of the black hole being non/rotating, is non zero due to the choice of the coordinates. The remaining panels, with 1D profiles along the $z$ axis, show the numerical error of three representative quantities at the final time $t=1000$ with respect to the initial equilibrium solution.